Oxygen (O2) gas is observed to diffuse through a capillary for 18 seconds. How long will it take for methane gas (CH4) to diffuse through the same orifice under identical conditions of temperature and pressure? (10 points)


1
Expert's answer
2022-01-19T10:36:03-0500

R2/R1 = M1/M2

M (O2) = 32 g/mol

M (CH4) = 16 g/mol

R2/18 = 32/16

R2(CH4) = 32/16 x 18 = 36 seconds